### v2

@@ 5 USE CASES
The utilization of ORAN-based solutions for NTNs will enable a variety of innovative use cases in the upcoming 6G networks. In this section, we present three representative use cases for the proposed AI-Native ORAN-based NTN framework, including emergency communication, remote areas coverage, and V2X communications, which exploit its advantages of fast deployment, cost efficiency, and edge computing capability.
5.1 EMERGENCY COMMUNICATION
@@
The emergency communication requires the fast deployment of networks to replace the destroyed terrestrial infrastructure and provide additional capacity to support post-disaster recovery activities. Although NTN has been regarded as a promising solution for emergency communication due to its large coverage, the fixed orbit of NTN limits its application to provide timely connectivity in emergency communication. For example, in order to provide additional capacity to the disaster area, the satellite orbit transfer techniques are utilized to schedule the satellites and change their orbital parameters, such as altitude, inclination, or eccentricity, which may take several days for a single MNO to establish end-to-end links via the Hohmann transfer orbit. However, the proposed ORAN-based NTN architecture enables interoperation between MNOs, allowing multiple MNOs to coordinate their access link, FH link, midhaul link, and backhaul link resources to promptly establish the emergency communication network. Furthermore, ORAN enables network slicing, allowing for the creation of dedicated, prioritized network slices for emergency services. This ensures that critical communications, such as voice, data, drone video feeds, and sensor information, are allocated sufficient bandwidth, even during network congestion. Additionally, Near-RT and Non-RT RICs dynamically and swiftly allocate resources, further prioritizing emergency communications and facilitating effective crisis management.
5.2 REMOTE AREAS COVERAGE
@@
Although NTNs have the potential to provide ubiquitous connections, a single MNO may not be willing to invest in sufficient NTN infrastructure for remote areas due to the high cost of testing, deployment, and operation. In LEO satellite-based NTNs, a single operator may need to deploy anywhere from several hundred to thousands of satellites to establish its own network, where each satellite provides service only while it is within the range of a particular geographic area. Given the lower penetration and shorter propagation distances of mmWave and sub-THz waves, a significantly higher number of BSs will be required for deployment in NTN. To lower the substantial deployment cost associated with such a system, operators across different countries can benefit from sharing LEO satellite infrastructure by standardizing the interfaces between ground stations and satellites, and between satellites. This is the same concept as sharing O-RU/O-DU among operators in ORAN. Therefore, ORAN-based NTN has the potential to significantly reduce the cost of providing connections to remote areas, based on on-orbit test and upgrade advantages along with the collaboration of multiple MNOs.
5.3 V2X COMMUNICATIONS
In B5G and 6G networks, V2X communications are expected to advance substantially, enabling applications such as autonomous driving, cooperative awareness, and real-time traffic coordination. The 5G Automotive Association (5GAA) has emphasized the potential of NTNs for connected vehicles, with early deployments anticipated by 2027 and broader adoption by 2030 [5gaa2024ntn]. The proposed ORAN-based NTN architecture can effectively address key V2X challenges by combining the flexibility of ORAN with the wide-area coverage of satellites and HAPs. This integration ensures reliable connectivity in underserved remote areas, which is essential for remote driving support functions such as safety alerts and collision avoidance [oran2025private].
@@ 5.3 V2X COMMUNICATIONS
Furthermore, the ultra-low latency can be supported for safety-critical V2X communication by leveraging edge computing at NTN platforms. The AI-enabled RICs in the proposed framework help dynamically optimize vehicular communication, prioritize traffic during congestion or emergencies, and adapt to vehicle mobility in real time. Additionally, ORAN’s programmable network slicing can extend over NTN links to deliver dedicated slices for different vehicular needs, ensure URLLC for autonomous control, and support enhanced mobile broadband (eMBB) slices for diagnostics [alam2024slicing]. This ability to tailor services across hybrid terrestrial and non-terrestrial network, will be critical for supporting future heterogeneous V2X demands.
@@ 6 FUTURE RESEARCH DIRECTIONS
Building upon the state-of-the-art literature review, the proposed ORAN-based NTN framework, and the representative use cases presented above, we further conduct an in-depth analysis of prospective future research directions as shown in Fig. 8.
@@
6.1 INFRASTRUCTURE SHARING
6.1.1 DEFINITION
@@
Infrastructure sharing among mobile network operators is a key business model in the industry, with competitors turning into partners to minimize the investment costs and expand the coverage of MNOs [GSMA2025]. In a traditional TN, infrastructure sharing can be broadly classified into two categories: passive and active. Passive sharing involves the sharing of telecommunication sites, towers, buildings, and facilities such as power supplies and air conditioning. Active sharing, on the other hand, involves the components of the active network layer, such as the antennas, base stations, or even core network elements [ITU2008_sharingInfrastructure]. Due to the increasing development and deployment costs in NTN, the infrastructure sharing in the NTN environment remains an urgent challenge to be solved.
6.1.2 RECENT ADVANCES AND PROMISING DIRECTIONS
@@
To the best of our knowledge, few works have explicitly investigated or systematically analyzed NTN-based infrastructure sharing strategies [lee2025ran, he2025accountability]. In [lee2025ran], the authors proposed a joint power allocation and FH compression optimization in ORAN-based NTN for multiple MNOs sharing the LEO satellites, but it lacks policy-level mechanisms for sharing resources, such as spectrum, beams and computing resources, across multiple operators or verticals in NTN environments. In [he2025accountability], the authors analyzed several ORAN use cases related to collaboration and resource sharing among multiple parties, with little attention to NTN integration. Therefore, we present a brief discussion of NTN-based infrastructure sharing enabled by our proposed ORAN-based NTN framework, with the goal of offering insights and guidance for future studies in this area.
@@
In NTN-based infrastructure sharing, the active sharing will involve the sharing of the RAN, TPN, and CN, while passive sharing will also include the sharing of orbits or NTN platforms. Based on the proposed ORAN-based NTN framework, the active sharing for multiple NTN MNOs can be achieved based on the interoperability and NFV characteristics of ORAN. The interoperability allows each MNO to utilize other MNOs’ DUs or CUs to establish an E2E communication link under a dynamic network topology. Meanwhile, based on NFV, virtualized DU and CU pools can support the resource sharing among O-RUs and O-DUs of multiple MNOs [he2025accountability]. However, unlike terrestrial passive sharing, NTN-based passive sharing depends on the mobility and coverage characteristics of different platforms and faces several unique challenges. Specifically, UAVs and HAPs have limited regional coverage, which requires domestic coordination among MNOs. Satellites offer global coverage but demand cross-border or international cooperation, which is hindered by regulatory, privacy, and security concerns. Future research should focus on developing secure and efficient mechanisms for both domestic and international infrastructure sharing, as well as exploring the integration of privacy-preserving technologies such as blockchain.
6.2 HIERARCHICAL ML-BASED OPTIMIZATION
6.2.1 DEFINITION
@@
Hierarchical ML-based optimization refers to a multi-level learning and decision-making framework that decomposes complex network control tasks across different timescales or functional layers. Within the ORAN architecture, intelligent control is inherently structured around the Non-RT RIC and Near-RT RIC, naturally aligning with a hierarchical optimization framework across different timescales [oranwhitepaper-1]. The effectiveness of hierarchical ML in optimizing terrestrial ORAN systems has been extensively demonstrated in prior studies, such as traffic steering [Habib2025], network slicing [Ghafouri2024], resource allocation [You2023], etc.
6.2.2 RECENT ADVANCES AND PROMISING DIRECTIONS
@@
Several studies have explored hierarchical optimization frameworks in NTN for different objectives [shinde2024hierarchical, farajzadeh2025federated, umer2025intelligent]. The authors in [farajzadeh2025federated] proposed a hierarchical federated learning (HFL) framework for NTNs, where HAPS constellations act as intermediate servers to enable global-scale training. [shinde2024hierarchical] proposed a hierarchical reinforcement learning approach for joint network selection and computation offloading to achieve efficient resource utilization. [umer2025intelligent] developed a hierarchical deep reinforcement learning (HDRL) framework for intelligent spectrum sharing in integrated TN-NTN networks, enhancing both spectral efficiency and network capacity. However, these proposed frameworks remain decoupled from the ORAN architecture.
@@
In the NTN scenario, propagation delays vary significantly across NTN platforms due to variations in altitude, orbital dynamics, and processing capabilities limited by payload hardware and feeder links. Traditional hierarchical optimization mechanisms, which typically assume uniform or bounded latency, are therefore insufficient. Therefore, an important future direction is to develop ORAN-based hierarchical ML solutions that dynamically distribute intelligence across edge and cloud layers, ensuring efficient learning and decision-making under the stringent latency and connectivity constraints of NTNs.
6.3 MOBILITY MANAGEMENT
6.3.1 DEFINITION
Mobility management is a critical challenge in NTNs, mainly encompassing handover, location management, and mobility robustness optimization. In TN, handover decisions are primarily coverage-driven and are typically triggered by the degradation of signal strength indicators such as RSRP and RSRQ. As the user equipment (UE) approaches the cell edge, the serving BS relies on UE measurement reports and predefined handover strategies to determine whether a handover should be initiated [haghrah2023survey].
However, in the satellite-based NTN, mobility is predominantly driven by the movement of satellite beams rather than the UE’s motion. For LEO satellites orbiting at 7.8 km/s, beams sweep rapidly across the Earth’s surface, forcing even stationary users to experience frequent handovers every few minutes. This high handover rate, combined with large Doppler shifts, leads to increased UE power consumption and imposes considerable signaling and processing overhead on the network [alsaeedy2024survey]. Furthermore, the relatively constant UE-satellite distance within a beam footprint leads to minimal signal variations, making conventional measurement-based handover criteria less effective. Furthermore, inherent NTN characteristics such as long propagation delays, dynamic beam patterns, and intermittent feeder links may lead to outdated measurements and handover failures.
6.3.2 RECENT ADVANCES AND PROMISING DIRECTIONS
Recent studies have focused on analyzing and improving mobility management in NTNs to ensure reliable connectivity [alsaeedy2024survey, seeram2025handover, pasumarthy2026designing, 3gpp38821]. Considering the different payload deployment options in satellite-based NTN and the ORAN architecture, the handover procedures have been categorized into intra-DU (intra-satellite), inter-DU (inter-satellite), and inter-CU (inter-access) handovers in [pasumarthy2026designing]. 3GPP has specified several NTN mobility management trigger mechanisms, based on measurement, location, time, timing advance, and elevation angle, respectively [3gpp38821]. In [alsaeedy2024survey], the authors discussed different mobility management solutions for satellite-based 5G NTNs and evaluated their impacts on UE power consumption and signaling overhead, highlighting the potential of AI-based methods for dynamic mobility management optimization. The authors in [seeram2025handover] focused on optimizing handover performance in NTN to enhance UE effective service time, and also investigated the impact of different ORAN functional splits between ground stations and LEO satellites on handover delays. This work revealed that the highest availability was achieved with the gNB onboard the satellite.
The above studies have advanced mobility management in NTNs from different perspectives. To achieve more effective mobility management within the proposed AI-Native ORAN-based NTN architecture, it is important to consider developing ephemeris-aware xApps for proactive beam and satellite handovers, designing Near-RT RIC policies to mitigate Doppler and delay effects, and incorporating AI/ML models to capture the spatio-temporal evolution of beam footprints [baranda2025architectural]. In addition, lightweight inference frameworks are required to ensure reliable decision-making under backhaul constraints. Furthermore, the multi-constellation and hybrid mobility mechanisms should be explored for seamless connectivity. Finally, integrating operator-defined policies into the RIC will be essential to align mobility management with both network performance and service-level objectives.
6.4 EDGE AI
6.4.1 DEFINITION
@@
Edge AI typically refers to the integration of AI and ML capabilities at the network edge, enabling real-time decision-making and intelligent automation, and reducing the need for extensive data transmission [letaief2021edge]. The envisioned solution of integrated NTN and TN in future 6G networks will require the edge-AI-enabled network to support the desired use cases, such as autonomous driving, V2X using vehicular edge computing, virtual reality (VR), and their stringent reliability requirements [AI-for-TN-NTN-1].
6.4.2 RECENT ADVANCES AND PROMISING DIRECTIONS
@@
Recent studies have further explored the integration of Edge AI into ORAN and NTN systems to enable the onboard decision making [firouzi20252] and enhance the physical layer security [chou2024edgeai]. The importance and the deployment design of Edge AI in the NTN scenario are discussed in [garcia2026edge]. In ORAN-based NTNs, future research should focus on efficiently partitioning AI tasks between edge and cloud, by keeping latency-sensitive inference on NTN platforms, while offloading computation-heavy training to ground controllers.
@@
Large language models (LLMs) have recently emerged as a powerful tool to improve the generalizability and adaptability of AI-driven network management [salmi2025ai]. Recent studies have demonstrated the feasibility of integrating LLMs into the ORAN architecture. The authors in [tang2025end] introduced an edge AI orchestration framework that leverages LLM-based rApps for interactive and adaptive network service management. [wu2025llm] proposed an ORAN system deploying LLM-empowered xApps in the Near-RT RIC to enable resilient, real-time decision-making. Traditional machine learning models, which rely primarily on offline training, often struggle to adapt to the highly dynamic and heterogeneous conditions of NTN environments. In contrast, LLM-based agents possess strong reasoning and contextual understanding capabilities, allowing them to respond adaptively to changing network states. Given the inherent mobility and temporal variability of NTN environments, future research may focus on domain-specific fine-tuning of LLM, and collaborative reasoning among distributed LLM agents in ORAN-based NTN, to enable intelligent, robust, and self-evolving network management.
6.5 CROSS-DOMAIN AI
6.5.1 DEFINITION

### v3

@@ 5 USE CASES
The utilization of ORAN-based solutions for NTNs will enable a variety of innovative use cases in the upcoming 6G networks. In this section, we present three representative use cases of the proposed AI-Native ORAN-based NTN framework, including emergency communication, remote areas coverage, and V2X communications. The realization of these use cases is fundamentally supported by the proposed framework through fast deployment, better cost efficiency, and reliable connectivity, as shown in Fig. 9.
Figure 9: Use cases of the AI-Native ORAN-based NTN framework.
5-A EMERGENCY COMMUNICATION
@@
Emergency communication usually requires the fast deployment of networks to replace destroyed terrestrial infrastructure and provide additional capacity for post-disaster recovery. Although the satellites are promising for such scenarios due to their wide coverage, the fixed orbit limits their ability to offer timely connectivity. Satellite orbit transfer techniques can be used to adjust satellite orbits by configuring orbital parameters such as altitude, inclination, and eccentricity. However, for a single MNO, establishing E2E links via the Hohmann transfer orbit may take several days to complete. To overcome these limitations, the proposed ORAN-based NTN framework supports interoperation among MNOs, allowing multiple MNOs to coordinate and share the access, fronthaul, midhaul, and backhaul link resources. Through the multi-MNO collaboration, an emergency communication network can be established more rapidly, without relying solely on slowly adjusting the satellites’ orbit from a single MNO. Furthermore, since ORAN can support network slicing, the proposed framework can create dedicated and prioritized network slices for emergency services. This ensures that critical communications, such as voice, data, drone video feeds, and sensor information, are allocated sufficient bandwidth even under network congestion. In addition, the RICs can dynamically allocate resources to further prioritize emergency communication traffic, thereby supporting timely and effective crisis management.
5-B REMOTE AREAS COVERAGE
@@
Although NTNs offer the potential for ubiquitous connectivity, the high costs associated with testing, deployment, and operation may discourage a single MNO from investing in sufficient NTN infrastructure for remote areas coverage. For example, in LEO satellite-based NTNs, a single MNO would need to deploy hundreds to thousands of satellites to build its own network, with each satellite serving a limited geographic area over a limited time window. Moreover, to reduce the substantial deployment cost of base stations in NTN, MNOs across different regions can benefit from sharing the NTN infrastructure. This can be realized through the standardization of interfaces between base stations and NTN platforms, as well as among NTN platforms operated by different MNOs. This approach follows the same principle as O-RU and O-DU sharing in ORAN. Based on the on-orbit test and upgrade, along with the collaboration of multiple MNOs, the proposed ORAN-based NTN framework has the potential to significantly reduce the costs of remote areas coverage.
5-C V2X COMMUNICATIONS
@@
In B5G and 6G networks, V2X communications are expected to advance substantially, enabling applications such as autonomous driving, cooperative awareness, and real-time traffic coordination. The 5G Automotive Association (5GAA) has emphasized the potential of NTNs for connected vehicles, with early deployments anticipated by 2027 and broader adoption by 2030 [7]. By combining the virtualization and intelligence of ORAN with the flexibility and wide coverage of NTN, the proposed ORAN-based NTN framework can effectively address the existing key challenges in V2X communication. Specifically, the ORAN-based NTN framework can provide reliable connectivity in underserved remote areas, which makes it also suitable for supporting remote driving functions, such as safety alerts and collision avoidance [95]. Moreover, the ultra-low latency can be supported for safety-critical V2X communication by leveraging edge computing at NTN platforms. The AI-enabled RICs in the proposed framework help dynamically optimize vehicular communication, prioritize traffic during congestion or emergencies, and adapt to vehicle mobility in real time. Additionally, ORAN’s programmable network slicing can extend over NTN links to deliver dedicated slices for different vehicular needs, ensure URLLC for autonomous control, and support enhanced mobile broadband (eMBB) slices for diagnostics [14]. This capability to tailor services across hybrid terrestrial and non-terrestrial networks will be important for supporting future heterogeneous V2X demands.
@@ 6 FUTURE RESEARCH DIRECTIONS
In this section, we further conduct an in-depth analysis of the prospective future research directions for the proposed framework, which is shown in Fig. 10. These directions encompass both business and technical aspects of AI-Native ORAN-based NTNs. Specifically, the business perspective focuses on infrastructure sharing, which enables cost-efficient and collaborative network deployment. From the technical perspective, several promising AI/ML methods are identified, including hierarchical ML-based optimization, edge AI, and cross-domain AI. Furthermore, at the system level, mobility management and digital twin are investigated, serving as essential mechanisms for effective network control and accurate predictive system modeling.
@@
6-A INFRASTRUCTURE SHARING
6-A.1 DEFINITION
@@
Infrastructure sharing among MNOs is a key business model in the industry, with competitors turning into partners to minimize the investment costs and expand the coverage of MNOs [51]. In a traditional TN, infrastructure sharing can be broadly classified into two categories: passive and active. Passive sharing involves the sharing of telecommunication sites, towers, buildings, and facilities such as power supplies and air conditioning. Active sharing, on the other hand, involves the components of the active network layer, such as the antennas, base stations, or even core network elements [62]. Due to the increasing development and deployment costs in NTN, the infrastructure sharing in the NTN environment remains an urgent challenge to be solved.
6-A.2 RECENT ADVANCES AND PROMISING DIRECTIONS
@@
To the best of our knowledge, few works have explicitly investigated or systematically analyzed NTN-based infrastructure sharing strategies [71, 56]. In [71], the authors proposed a joint power allocation and FH compression optimization in ORAN-based NTN for multiple MNOs sharing the LEO satellites, but it lacks policy-level mechanisms for sharing resources, such as spectrum, beams, and computing resources, across multiple operators or verticals in NTN environments. In [56], the authors analyzed several ORAN use cases related to collaboration and resource sharing among multiple parties, with little attention to NTN integration. Therefore, we provide a brief discussion on NTN-based infrastructure sharing under the proposed ORAN-based NTN framework, aiming to offer preliminary insights and serve as a reference for future studies in this area.
@@
In NTN-based infrastructure sharing, the active sharing will involve the sharing of the RAN, TPN, and CN, while passive sharing will also include the sharing of orbits or NTN platforms. Based on the proposed framework, the active sharing for multiple NTN MNOs can be achieved based on the interoperability and NFV of ORAN. The interoperability allows each MNO to utilize other MNOs’ DUs or CUs to establish an E2E communication link under a dynamic network topology. NFV enables virtualized DU and CU pools, which can support the resource sharing among O-RUs and O-DUs of multiple MNOs [56]. However, unlike terrestrial passive sharing, NTN-based passive sharing depends on the mobility and coverage characteristics of different NTN platforms and faces several unique challenges. Specifically, UAVs and HAPs have limited regional coverage, which requires domestic coordination among MNOs. Satellites offer global coverage but demand cross-border or international cooperation, which is hindered by regulatory, privacy, and security concerns. Therefore, considering the above challenges, future research can focus on developing secure and efficient mechanisms for both domestic and international infrastructure sharing, as well as exploring the integration of privacy-preserving technologies such as blockchain.
6-B HIERARCHICAL ML-BASED OPTIMIZATION
6-B.1 DEFINITION
@@
Hierarchical ML-based optimization refers to a multi-level learning and decision-making framework that decomposes complicated network control tasks across different timescales or functional layers. Within the ORAN architecture, intelligent control is inherently structured around the Non-RT RIC and Near-RT RIC, which naturally aligns with the hierarchical optimization framework across different timescales [92]. Prior studies have demonstrated the effectiveness of hierarchical ML for optimizing terrestrial ORAN systems in areas such as traffic steering [52], network slicing [49], and resource allocation [132].
6-B.2 RECENT ADVANCES AND PROMISING DIRECTIONS
@@
Several studies have explored the hierarchical ML-based optimization framework in NTN for different objectives [114, 46, 121]. The authors in [46] proposed a hierarchical federated learning (HFL) framework for NTNs, where HAPs constellations act as intermediate servers to enable global-scale training. [114] proposed a hierarchical reinforcement learning approach for joint network selection and computation offloading to achieve efficient resource utilization. [121] developed a hierarchical deep reinforcement learning (HDRL) framework for intelligent spectrum sharing in integrated TN-NTN networks, enhancing both spectral efficiency and network capacity. However, these proposed frameworks remain decoupled from the ORAN architecture.
@@
In NTN scenarios, propagation delays can vary significantly across different platforms due to differences in altitude and orbital dynamics. In addition, processing capabilities are constrained by onboard payload hardware and the limited feeder link. Traditional hierarchical optimization mechanisms, which typically assume uniform or bounded latency, are therefore insufficient. Therefore, an important future direction is to develop ORAN-based hierarchical ML solutions that can dynamically distribute intelligence across edge and cloud layers, and ensure efficient learning under the stringent latency constraints of NTNs.
6-C EDGE AI
6-C.1 DEFINITION
@@
Edge AI typically refers to the integration of AI and ML capabilities at the network edge, enabling real-time decision-making and reducing the need for extensive data transmission [72]. In future 6G networks, the integration of NTN and TN will rely on edge AI to support the emerging use cases that have stringent reliability requirements, such as autonomous driving, vehicular edge computing in V2X network, and virtual reality (VR) [28].
6-C.2 RECENT ADVANCES AND PROMISING DIRECTIONS
@@
Recent studies have further explored the integration of edge AI into ORAN and NTN systems to enable the onboard decision making [47] and enhance the physical layer security [31]. The importance and the deployment design of edge AI in the NTN scenario are discussed in [48]. In ORAN-based NTNs, one of the important future directions is to achieve intelligent partitioning of AI tasks between edge and cloud. This requires keeping latency-sensitive inference on NTN platforms, while offloading computation-heavy training to ground controllers. Meanwhile, since NTN platforms have stringent and different power constraints, another future direction is to focus on enhancing the energy efficiency of AI-driven RAN functions and control loops, which is essential for ensuring long-duration and cost-effective NTN operation.
@@
In addition, LLMs have recently emerged as a powerful tool to improve the generalizability and adaptability of AI-driven network management [106]. Recent studies have demonstrated the feasibility of integrating LLMs into the ORAN architecture or edge AI. The authors in [118] introduced an edge AI orchestration framework that leverages LLM-based rApps for interactive and adaptive network service management. [131] proposed an ORAN system deploying LLM-empowered xApps in the Near-RT RIC to enable resilient, real-time decision-making. Traditional machine learning models, which rely primarily on offline training, often struggle to adapt to the highly dynamic and heterogeneous conditions of NTN environments. In contrast, LLM-based agents possess strong reasoning and contextual understanding capabilities, allowing them to respond adaptively to changing network conditions. Therefore, to achieve a more robust network management, future studies may focus on domain-specific fine-tuning of LLMs in rAPPs or xApps for NTN, as well as collaborative reasoning among distributed LLM agents in ORAN-based NTNs.
6-D CROSS-DOMAIN AI
6-D.1 DEFINITION